1. A 16.2 m long wooden log has uniform diameter of 2 m. to what length the log should be cut to obtain a piece of 22m3 Volume.
Answer by Cromlix(4381)

Volume of a cylinder = Pi r^2 h
22m^3 = Pi * 1^2 h
22m^3/Pi = height (length)
length = 7m
Hope this helps.
:-)